OpenStudy (anonymous):

Find GCD for a and b using Euclidean Algorithm.let a = 427, b = 154 d = sa + tb

OpenStudy (anonymous):

Here is my solutions: 427 = 2 (154 )+ 119 154 = 1 (119) + 35 119 = 3 (35) + 14 35 = 2 (14) + 7 14 = 2 (7) + 0 d = GCD (427, 154) = 7 d = sa + tb 35 = 2 (14) + 7 7 = 35 - 2 [14] 7 = 35 - 2 [119 - 3 [35]] 7 = 35 - 2 [119 - 3[154-119]] 7 = 35 - 2[119] + 6[154-119]] 7 = [154-119] - 2[119] + 6[154] - 6[119] 7 = [154-119] -8[119] + 6[154] 7 = 7[154] -9[427-2[154]] 7 = 7[154] -9[427] + 18[154] 7 = 25[154] - 9[427] By comparing: d = sa + tb a = -9 b = 25 Do anyone using more simple ways to find GCD method.